The sum of money amounting to $13.45 consist of nickels, dimes and quarters. There are three times as many quarters as there are dimes and there are 29 more nickels than quarters. How many of each coin?
Answer by mananth(12270)

dimes =x
quarters = 3x
nickels = 3x+29
10x+3*25x+5(3x+29)=1345
10x+75x+15x+145=1345
100x=1345-145
100x=1200
x=12 dimes
3x===>3*12=36 quarters
3x+29=3*12+29=65 nickels
CHECK
120+900+325=1345